Offwego Airlines has a daily flight from Chicago to Las Vegas. On average, 13 ticket holders cancel their reservations, so the company intentionally overbooks the flight. Cancellations can be described by a normal distribution with a mean of 13 passengers and a standard deviation of 5.30 passengers. Profit per passenger is $93. If a passenger arrives but cannot board due to overbooking, the company policy is to provide a cash payment of $193.
Use Table B.

How many tickets should be overbooked to maximize expected profit? (Round your z value to 2 decimal places. Round your other intermediate calculations to 4 decimal places and final answer to the nearest whole number.)
Answer #1

Cu = $93

Co = $193

Service level = 93 / (193 + 93) = 0.325174825

For 0.325174825 service level, Z = - 0.45

Overbook: 13 – 0.45 (5.3) = 10.615 or 11 tickets.